Assume that in the year 2007 stock and net fixed assets amounted to ksh 22 m and 30 m respectively sales amounted to 50 m. Calculate the stock and fixed assets as a percentage of sales.
1
Expert's answer
2018-11-20T15:48:10-0500

ROA=(net income)/(total assets)=23/1787=1.29%